# Boosting In-Context Learning in LLMs Through the Lens of Classical Supervised Learning

### v1

@@ Abstract
In-Context Learning (ICL) allows Large Language Models (LLM) to adapt to new tasks with just a few examples, but their predictions often suffer from systematic biases, leading to unstable performances in classification. While calibration techniques are proposed to mitigate these biases, we show that, in the logit space, many of these methods are equivalent to merely shifting the LLM’s decision boundary without having the ability to alter its orientation. This proves inadequate when biases cause the LLM to be severely misdirected. To address these limitations and provide a unifying framework, we propose Supervised Calibration (SC), a loss-minimization based framework, which learns an optimal, per-class affine transformation of LLM’s predictive probabilities in the logit space without requiring external data beyond the context. By using a more expressive functional class, SC not only subsumes many existing calibration methods in ICL as special cases but also enables the ability of altering and even completely reversing the orientation of the LLM’s decision boundary. Furthermore, SC’s loss-based nature facilitates the seamless integration of two purpose-built regularization techniques—context-invariance and directional trust-region regularizers. The former is designed to tackle the instability issue in ICL, while the latter is to control the degree of calibration. Finally, SC delivers state-of-the-art performance over calibration baselines in the 4-shot, 8-shot, and 16-shot settings across all nine datasets for Mistral-7B-Instruct-v0.3, Llama-2-7B-chat, and Qwen2-7B-Instruct.
@@ 1 Introduction
State-of-the-art LLMs exhibit a striking in-context learning (ICL) capability: with only a handful of input–label exemplars, they generalize to unseen queries almost as if they had been fine-tuned, thus functioning as highly sample-efficient few-shot learners [2, 13]. However, a growing body of evidence shows that ICL performance can be brittle—sometimes systematically biased—with respect to seemingly innocent design choices such as template wording [17], verbaliser selection [7], and the particular demonstrations given [12]. These biases and sensitivity of ICL pose a practical barrier to developing applications that are both adaptable and robust.
Motivated by this, extensive research has been conducted to develop calibration approaches to address such a challenge for classification problems in ICL. The majority of calibration methods fall under label-marginal-based calibration (LM) [42, 8, 5, 44]. These methods first estimate the LLM’s probability for each label given the context alone via various approaches. They then discount the predictive probabilities of the LLM for the labels that are over-represented and boost those that are under-represented. See detailed discussion in the later sections.
@@
Despite the empirical success of these methods, their ability of correcting the predictive probabilities of the LLM via its internal estimated prior is limited. Specifically, we show in Section 3.4 that the underlying idea of these methods is equivalent to optimally shifting the decision threshold of the base LLM. Hence, they are inherently incapable of altering or reversing the orientation of the decision boundary. This becomes problematic when the base LLM performs poorly. To further illustrate this limitation, consider a binary classification problem in Figure 1 (a), where the base LLM only achieves $30\%$ accuracy. Since LM methods can only shift the decision threshold, their maximum improvement over the base LLM is capped, only achieving the level of random guessing as seen in Figure 1 (b). One may expect that such an issue becomes more common and severe in the multiclass classification, where distinguishing among a larger number of labels is inherently more difficult. For instance, on the SST-5 dataset, the average accuracy across three representative LLMs is only 22%, highlighting the severity of this challenge. This limitation motivates the need for a more principled calibration framework—one that is capable of correcting severely misaligned LLM predictions when necessary (e.g., by reversing the decision direction), and that subsumes existing methods as special cases while remaining both theoretically grounded and practically robust.
@@
To achieve this goal, we introduce Supervised Calibration (SC), which is motivated by conceptualizing existing approaches as learning a calibrated classifier: they take a LLM’s logits as input features and subsequently optimize a bias term to shift these logits. However, this shift only corresponds to moving the LLM’s decision boundary to maximize the predictive accuracy illustrated in Figure 1 (b). Therefore, to enable more comprehensive adjustments—specifically, the ability to alter or reverse the orientation of the LLM’s decision boundary–the proposed SC leverages the paradigm of loss-function-based classification and optimize both the bias and the scaling factor jointly. Our approach begins by generating a surrogate dataset, removing the necessity of external dataset beyond the given context. From this surrogate data, we extract features in the form of logits derived from the base LLM’s output probabilities. Then we employ these features, paired with their corresponding true labels to train a standard classifier, which learns not only an optimal bias term but also an optimal rescaling factor. Critically, the concurrent optimization of this rescaling factor empowers our approach to reverse the LLM’s decision boundary when advantageous (as illustrated in Figure 1 (c)). Moreover, the loss-minimization framework underpinning SC inherently supports the integration of regularization techniques designed for addressing the common problems in ICL and calibration. In this context, we propose a novel context-invariance regularizer for addressing the instability issue in ICL and a directional trust-region regularizer for controlling the degree of calibration. From a statistical viewpoint, these characteristics allow SC to pursue a balance regarding to the bias-variance trade-off. While SC’s flexibility targets a reduction in approximation error over LM methods, its regularization components actively constrain variance—an essential consideration within the data-scarce ICL paradigm. Collectively, SC delivers an adaptable, stable, and theoretically grounded framework that improves LLMs’ classification quality in few-shot settings, enabling fairer and more socially impactful applications as a result. Experimental results demonstrate that SC consistently outperforms existing calibration methods across a broad range of tasks, significantly enhancing the predictive performance of three distinct LLMs evaluated on nine inference datasets. For example, the performance of SC is striking on the SST-5 dataset with the Qwen model (8-shot setting), where it significantly outperforms baseline methods with accuracy from  25% (baselines) to 44%. This notable boost is directly attributable to its learned negative scaling factor which re-orients the base LLM decision boundary in this multiclass classification task. See Figure 4 for more details.
@@
Our main contributions are summarized as follows: Firstly, we propose Supervised Calibration, which adopts loss minimization framework from classical supervised learning and calibrates ICL via learning optimal bias and scaling factors, enabling not only shifting but also altering the orientation of the base LLM decision boundary; Secondly, we integrate the context-invariance and directional trust region regularizations in SC, enhancing the stability of ICL and controlling the degree of the calibration respectively; Thirdly, we provide a theoretical intuition behind SC and its generalization over the LM methods; Lastly, we conduct extensive empirical studies to demonstrate the state-of-the-art performance of SC over several existing calibration baselines. We have released the code to facilitate reproducibility. [^1] The code repository: https://github.com/gundemkorel/ICL
@@ 2 Related work
Diagnosing biases and calibration via Label Marginal A seminal study by Zhao et al. [42] identified three primary ICL biases—majority-label bias, recency bias, and common-token bias—and introduced Contextual Calibration (CC), a method that adjusts conditional probabilities by normalizing against outputs from content-free prompts. Subsequently, Holtzman et al. [8] observed that superficially distinct responses compete for probability mass, resulting in a degraded performance. Therefore, they proposed DCPMI, which recalibrates logits based on domain-conditional pointwise mutual information. Recent studies have uncovered additional sources of instability in ICL, including feature bias [27], positional biases within multiple-choice prompts [33, 21, 43], and domain-label biases [5]. Each study paired the diagnosis of a specific bias with a corresponding lightweight calibration strategy. For instance, Domain-Context Calibration (DC) proposed correcting predictions by averaging label distributions over random in-domain strings [5]. More recently, Zhou et al. [44] proposed Batch Calibration (BC), which estimates a contextual prior from unlabeled inference-time mini-batches and adjusting each example’s prediction probability accordingly. While these LM methods show empirical improvements, they are susceptible to failure when the base LLM substantially misaligned with the downstream task as they are unable to alter the LLM’s decision direction. This motivates exploring calibration frameworks with greater flexibility.
Calibration via centroids A parallel line of work mitigates in-context biases by replacing the standard decision rule with centroid-based classification. Han et al. [6] proposed Prototypical Calibration, which models output probability vectors using Gaussian mixtures and assigns labels based on cluster likelihood, improving robustness to prompt variation and class imbalance. Similarly, Cho et al. [3] introduced Hidden Calibration, which operates in the model’s latent space by computing class centroids over hidden states and classifying based on proximity. Although these methods show empirical performance gains, they rely on additional data beyond the in-context examples, which may not always be available or compatible with the ICL setting.
@@ 3.1 Background
Consider an $n$-class classification task with label verbaliser set $\mathcal{Y}=\{y_{0},\dots,y_{n-1}\}$ and query space $\mathcal{X}$. In few-shot in-context learning (ICL), the context $C_{k}$ is constructed by concatenating $k$ input–label exemplars $(x^{(i)},y^{(i)})$ formatted via a template function $T$ such that $C_{k}=\text{Concat}(T(x^{(1)},y^{(1)}),\dots,T(x^{(k)},y^{(k)}))$. Then given the context of $k$-shots and a testing query $x\in\mathcal{X}$, the LLM predicts a label via computing
@@
While ICL offers an appealing alternative to the gradient-based fine-tuning by allowing LLMs to adapt to new tasks via only a handful of in-prompt demonstrations, the resulting posterior distribution $P_{\text{LLM}}(y\mid x,C_{k})$ is often distorted by some systematic biases. Such biases inherent in ICL often stems from context examples, their order, formatting, or verbalizer choices, which makes $P_{\text{LLM}}(y\mid x,C_{k})$ significantly diverge from ground-truth posterior $P^{*}(y|x)$. Therefore, the objective of calibration is to refine LLM’s predictive probabilities $P_{\text{LLM}}(\cdot\mid x,C_{k})$ to aligns with $P^{*}(y|x)$.
Existing approaches are mainly focused on correcting the prior distribution of the label via estimating the LLM’s internal prior given the context. Despite their successes, one can show that these approaches boils down to merely shifting the LLM’s decision boundary, lacking the ability to alter an LLM’s orientation. This limitation turns out to be essential especially in multi-class classification, where an LLM can easily make persistent mistakes. See Figure 1. Therefore, to further reduce the biases and align with $P^{*}(y\mid x)$ in such cases of substantial misorientation, we develop a more principle calibration called Supervised Calibration introduced below.
@@ 3.2 Our proposal
[…]
$\mathbb{E}_{x\sim P^{\star}}\left[D_{\mathrm{KL}}\big{(}P^{\star}(\cdot\mid x)\,\|\,P_{\text{LLM}}(\cdot\mid x,C_{k})\big{)}\right],$
@@
$f^{\ast}$ $=\arg\min_{f\in\mathcal{F}}\mathbb{E}_{x\sim P^{\star}(x)}\left[D_{\mathrm{KL}}\big{(}P^{\star}(\cdot\mid x)\,\|\,f(P_{\mathrm{LLM}}(\cdot|x,C_{k}))\big{)}\right]$
$=\arg\min_{f\in\mathcal{F}}-\mathbb{E}_{(x,y)\sim P^{\star}}\left[\log(f_{y}(P_{\mathrm{LLM}}(\cdot|x,C_{k})))\right],$ (1)
where $f_{y}$ is the $y^{th}$-coordinate projection of $f$. Note that as long as $\mathcal{F}$ contains the identity map, applying $f^{\ast}$ enhances the fidelity of $P_{\text{LLM}}$.
To find $f^{\ast}$, we highlight two key challenges. Firstly, since our method is post-hoc, choosing an effective $\mathcal{F}$ operating solely on the base LLM predictive probabilities is essential. Secondly, there is no external data sampled from $P^{\ast}$ to approximate the objective function in Equation (1).
@@ 3.2.1 Affine-logit approximation and leave-subset-out strategy
Then, instead, we aim to choose the transformed function class ~=$\widetilde{\mathcal{F}}=\left\{f:\mathbb{R}^{n-1}\rightarrow\Delta^{n}\right\}$ for calibration. To facilitate it, notice that
$P^{\star}(y\mid x)$ $=\frac{P^{\star}(x\mid y)P^{\star}(y)}{P^{\star}(x)}\propto P_{\text{LLM}}(y\mid x,C_{k})\frac{P^{\star}(x\mid y)}{P_{\text{LLM}}(x\mid y,C_{k})}\;\frac{P^{\star}(y)}{P_{\text{LLM}}(y\mid C_{k})}$ (2)
@@
$L^{\star}_{c}(x)=m_{c}(x;C_{k})+\underbrace{\underbrace{\log\left(\frac{P^{\star}(x|c)P_{\text{LLM}}(x|0,C_{k})}{P^{\star}(x|0)P_{\text{LLM}}(x|c,C_{k})}\right)}_{\text{Class Conditional Shift}}+\underbrace{\log\left(\frac{P^{\star}(c)P_{\text{LLM}}(0|C_{k})}{P^{\star}(0)P_{\text{LLM}}(c|C_{k})}\right)}_{\text{Label Marginal Shift}}}_{\log\left(h(x,c,C_{k})/h(x,0,C_{k})\right)},$ (4)
@@
where $L^{\star}_{c}(x)=\log(P^{\ast}(c|x)/P^{\ast}(0|x))$ is the true logit for class $c$. Thus, the primary challenge of choosing $\mathcal{F}$ lies in approximating the unknown correction term $\log\left(h(x,c,C_{k})/h(x,0,C_{k})\right)$. Since we only have access to the LLM’s output logits $\mathbf{m}(x;C_{k})$, we propose to approximate $\{L^{\star}_{c}(x)\}_{c=1}^{n-1}$ via an affine transformation of $\{m_{c}(x;C_{k})\}_{c=1}^{n-1}$. In particular, our working model $L_{c}(x;\bm{\theta}_{c}^{k})$ is
$L_{c}(x;\bm{\theta}_{c}^{k})=w^{k}_{c}\,m_{c}(x;C_{k})+b^{k}_{c},\qquad c=1,\dots,n-1,$ (5)
@@
where $\bm{\theta}_{c}^{k}=(b^{k}_{c},w^{k}_{c})$ are calibration parameters associated with class $c$ and the context size $k$. This affine structure directly targets the two primary sources of discrepancies between true and LLM logits: class-conditional shift and label marginal shift as illustrated in Equation (4). Specifically, by rearranging Equation (5) as $L_{c}(x;\bm{\theta}_{c}^{k})=m_{c}(x;C_{k})+[(w^{k}_{c}-1)m_{c}(x;C_{k})+b^{k}_{c}]$, we see that the term $(w^{k}_{c}-1)m_{c}(x;C_{k})+b^{k}_{c}$ serves as our learned approximation to the true correction term $\log\left(h(x,c,C_{k})/h(x,0,C_{k})\right)$. Within this learned correction, the intercept $b_{c}^{k}$ primarily addresses the query-independent "Label Marginal Shift" component from Equation (4), compensating for discrepancies in label priors. The query-dependent term $(w_{c}^{k}-1)m_{c}(x;C_{k})$ targets the "Class Conditional Shift" by allowing the slope $w_{c}^{k}$ to rescale the LLM’s original logit $m_{c}(x;C_{k})$.
@@
Furthermore, $w_{c}^{k}$ enables the reorientation of the LLM’s decision boundary. For instance, a negative $w_{c}^{k}$ inverts the LLM’s initial assessment for a class relative to the reference, effectively correcting its predictive direction as illustrated in Figures 1 (c) and 4. This is a vital capability that methods merely learning a bias (i.e., fixing $w_{c}^{k}=1$) lack. As detailed in Section 3.4, our framework not only unifies but also generalizes several recent ICL calibration techniques. Finally, it naturally encompasses the base LLM’s original predictions as a special case when $b^{k}_{c}=0$ and $w^{k}_{c}=1$ for all $c$.
In terms of learning the parameters, if an external calibration dataset $\{(x^{(j)},y^{(j)})\}_{j=1}^{N_{cal}}$ is provided, we first compute the LLM’s logits $\mathbf{m}(x^{(j)};C_{k})$ for each $x^{(j)}$. Then based on Equation (1), we estimate the parameters via minimizing the negative log-likelihood, i.e.,
$\hat{\bm{\theta}}^{k}=\arg\min_{\bm{\theta}^{k}}\{\mathbb{L}_{k}(\bm{\theta}^{k})\triangleq-\sum_{j=1}^{N_{cal}}\log f_{y^{(j)}}(\mathbf{m}(x^{(j)};C_{k});\bm{\theta}^{k})\},$ (6)
where $\bm{\theta}^{k}=\{\bm{\theta}_{c}^{k}\}_{c=1}^{n-1}$ and $f_{c}(\mathbf{m}(x^{(j)};C_{k});\bm{\theta}^{k})=\frac{\mathbf{1}_{\{c>0\}}\exp(L_{c}(x;\bm{\theta}^{k}_{c}))+\mathbf{1}_{\{c=0\}}}{1+\sum_{i=1}^{n-1}\exp(L_{i}(x;\bm{\theta}^{k}_{i}))}$.
@@
This optimization problem is equivalent to standard multi-class logistic regression using the model logits $m_{c}$ as input features. However, there is no external calibration dataset available beyond $C_{k}$. Therefore, we propose generating surrogate training data directly from the demonstration context $C_{k}$ via a leave-subset-out strategy. Specifically, we first select a context size $i$ such that $i<k$. We then construct the surrogate training dataset $\mathcal{T}_{i}$ using Algorithm 1 in Appendix E, as illustrated in Figure 2. Finally, we estimate calibration parameters $\hat{\bm{\theta}}^{i}$ via minimizing $\mathbb{L}_{i}$ under $\mathcal{T}_{i}$. Note that this method can be applied across multiple context sizes $i$, enabling ensembling extensions of $\{\hat{\bm{\theta}}^{i}\}_{i\in I}$ to construct a final estimator for calibration.

### v3

@@ Abstract
In-Context Learning (ICL) allows Large Language Models (LLMs) to adapt to new tasks with just a few examples, but their predictions often suffer from systematic biases, leading to unstable performance in classification. While calibration techniques are proposed to mitigate these biases, we show that, in the logit space, many of these methods are equivalent to merely shifting the LLM’s decision boundary without having the ability to alter its orientation. This proves inadequate when biases cause the LLM to be severely misaligned. To address these limitations and provide a unifying framework, we propose Supervised Calibration (SC), a loss-minimization-based framework, which learns an optimal, per-class affine transformation of LLM’s predictive probabilities in the logit space without requiring external data beyond the context. By using a more expressive functional class, SC not only subsumes many existing calibration methods in ICL as special cases but also enables the ability of altering and even completely reversing the orientation of the LLM’s decision boundary. Furthermore, SC’s loss-based nature facilitates the seamless integration of two purpose-built regularization techniques, context-invariance and directional trust-region regularizers. The former is designed to tackle the instability issue in ICL, while the latter is to control the degree of calibration. Finally, SC delivers state-of-the-art performance over calibration baselines in the 4-shot, 8-shot, and 16-shot settings across all nine datasets for Mistral-7B-Instruct-v0.3, Llama-2-7B-chat, and Qwen2-7B-Instruct.
Machine Learning, ICML
@@ 1 Introduction
State‑of‑the‑art LLMs exhibit a striking in‑context learning (ICL) capability: with only a handful of input–label exemplars, they generalize to unseen queries almost as if they had been fine‑tuned, thus functioning as highly sample‑efficient few‑shot learners (Brown et al., 2020; Liu et al., 2021). However, a growing body of evidence shows that ICL performance can be brittle with respect to seemingly innocent design choices such as template wording (Min et al., 2022; Voronov et al., 2024) and the particular demonstrations given (Qin et al., 2024). These biases and sensitivities of ICL pose a practical barrier to developing applications that are both adaptable and robust. Motivated by this, extensive research has been conducted to develop calibration approaches to address such a challenge for classification problems in ICL. The majority of calibration methods fall under label-marginal-based calibration (LM). These methods first estimate the LLM’s probability for each label given the context alone via various approaches. They then discount the predictive probabilities of the LLM for the labels that are over-represented and boost those that are under-represented. See detailed discussion in the later sections.
@@
Despite the empirical success of these methods, their ability of correcting the predictive probabilities of the LLM via its internal estimated prior is limited. Specifically, we show in Section 3.4 that the underlying idea of these methods is equivalent to optimally shifting the decision threshold of the base LLM. Hence, they are inherently incapable of altering or reversing the orientation of the decision boundary. This becomes problematic when the base LLM performs poorly. To further illustrate this limitation, consider a binary classification problem in Figure 1 (a), where the base LLM only achieves $30\%$ accuracy. Since LM methods can only shift the decision threshold, their maximum improvement over the base LLM is capped, only achieving the level of random guessing as seen in Figure 1 (b). One may expect that such an issue becomes more common and severe in the multiclass classification, where distinguishing among a larger number of labels is inherently more difficult. For instance, on the SST-5 dataset, the average accuracy across three representative LLMs is only 22%, highlighting the severity of this challenge. This limitation motivates the need for a more principled calibration framework that is capable of correcting severely misaligned LLM predictions when necessary (e.g., by reversing the decision direction), and that subsumes existing methods as special cases while remaining both theoretically grounded and practically robust.
@@
To achieve this goal, we introduce Supervised Calibration (SC), which is motivated by conceptualizing existing approaches as learning a calibrated classifier: they take a LLM’s logits as input features and subsequently optimize a bias term to shift these logits. However, this shift only corresponds to moving the LLM’s decision boundary to maximize the predictive accuracy illustrated in Figure 1 (b). Therefore, to enable more comprehensive adjustments, specifically, the ability to alter or reverse the orientation of the LLM’s decision boundary, the proposed SC leverages the paradigm of loss-function-based classification and optimizes both the bias and the scaling factor jointly. Our approach begins by generating a surrogate dataset, removing the need for an external dataset beyond the given context. From this surrogate data, we extract features in the form of logits derived from the base LLM’s output probabilities. Then we employ these features, paired with their corresponding true labels to train a standard classifier, which learns not only an optimal bias term but also an optimal rescaling factor. Critically, the concurrent optimization of this rescaling factor empowers our approach to reverse the LLM’s decision boundary when advantageous (as illustrated in Figure 1 (c)). Moreover, the loss-minimization framework underpinning SC inherently supports the integration of regularization techniques designed for addressing the common problems in ICL and calibration. In this context, we propose a novel context-invariance regularizer for addressing the instability issue in ICL and a directional trust-region regularizer for controlling the degree of calibration. From a statistical viewpoint, these characteristics allow SC to pursue a balance with respect to the bias–variance trade-off. While SC’s flexibility targets a reduction in approximation error over LM methods, its regularization components actively constrain variance which is an essential consideration within the data-scarce ICL paradigm. Collectively, SC delivers an adaptable, stable, and theoretically grounded framework that improves LLMs’ classification quality in few-shot settings, potentially improving robustness across prompts and datasets as a result. Experimental results demonstrate that SC consistently outperforms existing calibration methods across a broad range of tasks, significantly enhancing the predictive performance of three distinct LLMs evaluated on nine inference datasets. For example, the performance of SC is striking on the SST-5 dataset with the Qwen model (8-shot setting), where it significantly outperforms baseline methods with accuracy from  25% (baselines) to 44%. This notable boost is directly attributable to its learned negative scaling factor which re-orients the base LLM decision boundary in this multiclass classification task. See Figure 4 for more details.
@@
Our main contributions are summarized as follows: Firstly, we propose Supervised Calibration, which adopts a loss minimization framework from classical supervised learning and calibrates ICL via learning optimal bias and scaling factors, enabling not only shifting but also altering the orientation of the base LLM decision boundary; Secondly, we integrate the context-invariance and directional trust-region regularizations in SC, enhancing the stability of ICL and controlling the degree of the calibration respectively; Thirdly, we provide a theoretical intuition behind SC and its generalization over the LM methods; Lastly, we conduct extensive empirical studies to demonstrate the state-of-the-art performance of SC over several existing baselines. [^1] The code for reproducibility: https://github.com/gundemkorel/ICL
@@ 2 Related Work
Diagnosing biases and calibration via Label Marginal. A seminal study by Zhao et al. (2021) identified primary in-context learning (ICL) biases—including majority-label, recency, and common-token bias—and introduced Contextual Calibration (CC), which adjusts probabilities by normalizing against content-free prompts. Subsequently, observing that competition for probability mass degrades performance, Holtzman et al. (2022) proposed DCPMI to recalibrate logits. Recent work has uncovered further ICL instabilities, such as feature and positional biases, with each diagnosis often paired with a lightweight calibration strategy (Si et al., 2023; Wang et al., 2023; Pezeshkpour and Hruschka, 2023). For instance, Domain-Context Calibration (DC) corrects predictions by averaging over random in-domain strings (Fei et al., 2023), while the more recent Batch Calibration (BC) uses unlabeled mini-batches to adjust each prediction (Zhou et al., 2023). Although these methods show empirical improvements, they can fail when the base LLM is substantially misaligned with the downstream task, as they cannot alter the model’s decision direction. This limitation motivates the exploration of calibration frameworks with greater flexibility.
Calibration via centroids. A parallel line of work mitigates in-context biases by replacing the standard decision rule with centroid-based classification. Han et al. (2022) proposed Prototypical Calibration, which models output probability vectors using Gaussian mixtures and assigns labels based on cluster likelihood, improving robustness to prompt variation and class imbalance. Similarly, Cho et al. (2024) introduced Hidden Calibration, which operates in the model’s latent space by computing class centroids over hidden states and classifying based on proximity. Although these methods show empirical performance gains, they rely on additional data beyond the in-context examples, which may not always be available or compatible with the ICL setting.
@@ 3.1 Background
Consider an nn‑class classification task with label verbaliser set $\mathcal{Y}=\{y_{0},\dots,y_{n-1}\}$ and query space $\mathcal{X}$. In few-shot in-context learning (ICL), the context $C_{k}$ is constructed by concatenating $k$ input–label exemplars $(x^{(i)},y^{(i)})$ formatted via a template function $T$ such that $C_{k}=\text{Concat}(T(x^{(1)},y^{(1)}),\dots,T(x^{(k)},y^{(k)}))$. Then given the context of $k$-shots and a testing query $x\in\mathcal{X}$, the LLM predicts a label via computing
@@
While ICL offers an appealing alternative to the gradient-based fine-tuning by allowing LLMs to adapt to new tasks via only a handful of in-prompt demonstrations, the resulting posterior distribution $P_{\text{LLM}}(y\mid x,C_{k})$ is often distorted by some systematic biases. Such biases inherent in ICL often stem from context examples or their order, which makes $P_{\text{LLM}}(y\mid x,C_{k})$ significantly diverge from ground-truth posterior $P^{*}(y|x)$. Therefore, the objective of calibration is to refine LLM’s predictive probabilities $P_{\text{LLM}}(\cdot\mid x,C_{k})$ to align with $P^{*}(y|x)$.
Existing approaches are mainly focused on correcting the prior distribution of the label via estimating the LLM’s internal prior given the context. Despite their successes, one can show that these approaches boil down to merely shifting the LLM’s decision boundary, lacking the ability to alter an LLM’s orientation. This limitation turns out to be essential especially in multi-class classification, where an LLM can easily make persistent mistakes. See Figure 1. Therefore, to further reduce the biases and align with $P^{*}(y\mid x)$ in such cases of substantial misorientation, we develop a more principled calibration called Supervised Calibration.
@@ 3.2 Our Proposal
To begin with, we assume the $k$ context examples $(x^{(i)},y^{(i)})_{i=1}^{k}\overset{\text{i.i.d.}}{\sim}P^{\ast}$. Due to the aforementioned biases, the LLM’s posterior $P_{\text{LLM}}(y\mid x,C_{k})$ can deviate notably from the truth $P^{*}(y\mid x)$. In particular, we measure their deviation via the Kullback–Leibler (KL) divergence defined as
$\mathbb{E}_{x\sim P^{*}}\left[D_{\mathrm{KL}}\big(P^{*}(\cdot\mid x)\,\|\,P_{\text{LLM}}(\cdot\mid x,C_{k})\big)\right],$
@@
$f^{\ast}$ $=\mathop{\arg\min}\limits_{\mathclap{f\in\mathcal{F}}}\text{ }\mathop{\mathbb{E}}\limits_{\mathclap{x\sim P^{*}}}[D_{\mathrm{KL}}(P^{*}(\cdot\mid x)\,\|\,f(P_{\mathrm{LLM}}(\cdot|x,C_{k})))]$
$=\mathop{\arg\min}\limits_{\mathclap{f\in\mathcal{F}}}\text{ }-\mathop{\mathbb{E}}\limits_{\mathclap{(x,y)\sim P^{*}}}[\log(f_{y}(P_{\mathrm{LLM}}(\cdot|x,C_{k})))],$ (1)
where $f_{y}$ is the $y^{th}$-coordinate projection of $f$. Note that as long as $\mathcal{F}$ contains the identity map, applying $f^{\ast}$ enhances the fidelity of $P_{\text{LLM}}$. To find $f^{\ast}$, we highlight two key challenges. Firstly, since our method is post-hoc, choosing an effective $\mathcal{F}$ operating solely on the base LLM predictive probabilities is essential. Secondly, there is no external data sampled from $P^{\ast}$ to approximate the objective function in Equation (1).
@@ 3.2.1 Affine-logit Approximation and Leave-subset-out Strategy
Then, instead, we aim to choose the transformed function class $\widetilde{\mathcal{F}}=\left\{f:\mathbb{R}^{n-1}\rightarrow\Delta^{n}\right\}$ for calibration. To facilitate it, notice that
$P^{*}(y\mid x)$ $=\frac{P^{*}(x\mid y)P^{*}(y)}{P^{*}(x)}$
$\propto P_{\text{LLM}}(y\mid x,C_{k})\frac{P^{*}(x\mid y)}{P_{\text{LLM}}(x\mid y,C_{k})}\;\frac{P^{*}(y)}{P_{\text{LLM}}(y\mid C_{k})}$
@@
$L^{\star}_{c}(x)$ $=m_{c}(x;C_{k})+\underbrace{\log\left(\frac{P^{*}(x|c)P_{\text{LLM}}(x|0,C_{k})}{P^{*}(x|0)P_{\text{LLM}}(x|c,C_{k})}\right)}_{\text{Class Conditional Shift}}$
$+\underbrace{\log\left(\frac{P^{*}(c)P_{\text{LLM}}(0|C_{k})}{P^{*}(0)P_{\text{LLM}}(c|C_{k})}\right)}_{\text{Label Marginal Shift}}$
$=m_{c}(x;C_{k})+\log\left(h(x,c,C_{k})/h(x,0,C_{k})\right)$ (3)
@@
where $L^{\star}_{c}(x)=\log(P^{\ast}(c|x)/P^{\ast}(0|x))$ is the true logit for class $c$. Thus, the primary challenge of choosing $\mathcal{F}$ lies in approximating the unknown correction term $\log\left(h(x,c,C_{k})/h(x,0,C_{k})\right)$ which is ”Class Conditional Shift” plus ”Label Marginal Shift”. Since we only have access to the LLM’s output logits $\mathbf{m}(x;C_{k})$, we propose to approximate $\{L^{\star}_{c}(x)\}_{c=1}^{n-1}$ via an affine transformation of $\{m_{c}(x;C_{k})\}_{c=1}^{n-1}$. In particular, our working model $L_{c}(x;\bm{\theta}_{c}^{k})$ for $c=1,\dots,n-1$ is
$L_{c}(x;\bm{\theta}_{c}^{k})=w^{k}_{c}\,m_{c}(x;C_{k})+b^{k}_{c},$ (4)
@@
where $\bm{\theta}_{c}^{k}=(b^{k}_{c},w^{k}_{c})$ are calibration parameters associated with class $c$ and the context size $k$. This affine structure directly targets the two primary sources of discrepancies between true and LLM logits: class-conditional shift and label marginal shift as illustrated in Equation (3.2.1). Specifically, by rearranging Equation (4) as
$L_{c}(x;\bm{\theta}_{c}^{k})=m_{c}(x;C_{k})+[(w^{k}_{c}-1)m_{c}(x;C_{k})+b^{k}_{c}],$
we see that the term $(w^{k}_{c}-1)m_{c}(x;C_{k})+b^{k}_{c}$ serves as our learned approximation to the true correction term $\log\left(h(x,c,C_{k})/h(x,0,C_{k})\right)$. Within this learned correction, the intercept $b_{c}^{k}$ primarily addresses the query-independent ”Label Marginal Shift” component from Equation (3.2.1), compensating for discrepancies in label priors. The query-dependent term $(w_{c}^{k}-1)m_{c}(x;C_{k})$ targets the ”Class Conditional Shift” by allowing the slope $w_{c}^{k}$ to rescale the LLM’s original logit $m_{c}(x;C_{k})$.
@@
Furthermore, $w_{c}^{k}$ enables the reorientation of the LLM’s decision boundary. For instance, a negative $w_{c}^{k}$ inverts the LLM’s initial assessment for a class relative to the reference, effectively correcting its predictive direction as illustrated in Figures 1 (c) and 4. This is a vital capability that methods merely learning a bias (i.e., fixing $w_{c}^{k}=1$) lack. As detailed in Section 3.4, our framework not only unifies but also generalizes several recent ICL calibration techniques. Finally, it naturally encompasses the base LLM’s original predictions as a special case when $b^{k}_{c}=0$ and $w^{k}_{c}=1$ for all $c$. In terms of learning the parameters, if an external calibration dataset $\{(x^{(j)},y^{(j)})\}_{j=1}^{N_{cal}}$ is provided, we first compute the LLM’s logits $\mathbf{m}(x^{(j)};C_{k})$ for each $x^{(j)}$. Then based on Equation (1), we estimate the parameters via minimizing the negative log-likelihood, i.e.,
$\hat{\bm{\theta}}^{k}$ $=\arg\min_{\bm{\theta}^{k}}\{\mathbb{L}_{k}(\bm{\theta}^{k})$
$\triangleq-\sum_{j=1}^{N_{cal}}\log f_{y^{(j)}}(\mathbf{m}(x^{(j)};C_{k});\bm{\theta}^{k})\},$ (5)
@@
where $\bm{\theta}^{k}=\{\bm{\theta}_{c}^{k}\}_{c=1}^{n-1}$ and $f_{c}(\mathbf{m}(x^{(j)};C_{k});\bm{\theta}^{k})=\frac{\mathbf{1}_{\{c>0\}}\exp(L_{c}(x;\bm{\theta}^{k}_{c}))+\mathbf{1}_{\{c=0\}}}{1+\sum_{i=1}^{n-1}\exp(L_{i}(x;\bm{\theta}^{k}_{i}))}$. This optimization problem is equivalent to standard multi-class logistic regression using the model logits $m_{c}$ as input features. However, there is no external calibration dataset available beyond $C_{k}$. Therefore, we propose generating surrogate training data directly from the demonstration context $C_{k}$ via a leave-subset-out strategy. Specifically, we first select a context size $i$ such that $i<k$. We then construct the surrogate training dataset $\mathcal{T}_{i}$ using Algorithm 1 in Appendix E, as illustrated in Figure 2. Finally, we estimate calibration parameters $\hat{\bm{\theta}}^{i}$ via minimizing $\mathbb{L}_{i}$ under $\mathcal{T}_{i}$. Note that this method can be applied across multiple context sizes $i$, enabling ensembling extensions of $\{\hat{\bm{\theta}}^{i}\}_{i\in I}$ to construct a final estimator for calibration.